# Enhancing Semi-Supervised Learning via Representative and Diverse Sample Selection

### v1

@@ Abstract
Semi-Supervised Learning (SSL) has become a preferred paradigm in many deep learning tasks, which reduces the need for human labor. Previous studies primarily focus on effectively utilising the labelled and unlabeled data to improve performance. However, we observe that how to select samples for labelling also significantly impacts performance, particularly under extremely low-budget settings. The sample selection task in SSL has been under-explored for a long time. To fill in this gap, we propose a Representative and Diverse Sample Selection approach (RDSS). By adopting a modified Frank-Wolfe algorithm to minimise a novel criterion $\alpha$-Maximum Mean Discrepancy ($\alpha$-MMD), RDSS samples a representative and diverse subset for annotation from the unlabeled data. We demonstrate that minimizing $\alpha$-MMD enhances the generalization ability of low-budget learning. Experimental results show that RDSS consistently improves the performance of several popular SSL frameworks and outperforms the state-of-the-art sample selection approaches used in Active Learning (AL) and Semi-Supervised Active Learning (SSAL), even with constrained annotation budgets.
@@ 2 Related Work
Semi-Supervised Learning. Semi-Supervised Learning (SSL) effectively utilizes sparse labeled data and abundant unlabeled data for model training. Consistency Regularization [33, 19, 42], Pseudo-Labeling [20, 53] and their hybrid strategies [37, 58, 34] are commonly used in SSL. Consistency Regularization ensures the model’s output stays stable even when there’s noise or small changes in the input, usually from the data augmentation [52]. Pseudo-labelling integrates high-confidence data pseudo-labels directly into training, adhering to entropy minimization [22]. Moreover, an integrative approach that combines the aforementioned strategies can also achieve substantial results [50, 55]. Even though these approaches have been proven effective, they usually assume that labelled samples are randomly selected from each class (i.e., stratified sampling), which is not practical in real-world scenarios where the label for each sample is unknown.
@@
Active Learning. Active learning (AL) aims to optimize the learning process by selecting the appropriate samples for labelling, reducing reliance on large labelled datasets. There are two different criteria for sample selection: uncertainty and representativeness. Uncertainty sampling selects samples about which the current model is most uncertain. Earlier studies utilized posterior probability [21, 46], entropy [17, 25], and classification margin [44] to estimate uncertainty. Recent research regards uncertainty as training loss [16, 56], influence on model performance [10, 23] or the prediction discrepancies between multiple classifiers [7]. However, uncertainty sampling methods may exhibit performance disparities across different models, leading researchers to focus on representativeness sampling, which aims to align the distribution of selected subset with that of the entire dataset [35, 36, 26]. Most AL approaches are difficult to perform well under extremely low-label settings. This may be because they usually require randomly selected samples to begin with and involve human annotators in iterative cycles of labelling and training, leading to substantial labelling overhead.
@@
Model-Free Subsampling. Subsampling is a statistical approach which selects a subset with size $m$ as a surrogate for the full dataset with size $n\gg m$. Model-free subsampling is preferred in data-driven modelling tasks, as it does not depend on the model assumptions. There are mainly two kinds of popular model-free subsampling methods. The one is induced by minimizing statistical discrepancies, which forces the distribution of subset to be similar to that of full data, in other words, selects representative subsamples, such as Wasserstein distance [12], energy distance [27], uniform design [59], maximum mean discrepancy [6] and generalized empirical $F$-discrepancy [60]. The other tends to select a diverse subset containing as many informative samples as possible [51]. The above-mentioned methodologies either exclusively focus on representativeness or diversity, which are difficult to effectively apply to SSL.
@@ 4 Representative and Diversity Sample Selection
The Representative and Diverse Sample Selection (RDSS) framework consists of two steps: (1) Quantification. We quantify the representativeness and diversity of selected samples by proposing a novel concept called $\alpha$-MMD (6), where $\lambda$ is replaced by $\alpha$ as the trade-off hyperparameter; (2) Optimization. We optimize $\alpha$-MMD by GKHR algorithm to obtain the optimally selected samples $\mathbf{X}_{\mathcal{I}^{*}_{m}}$.
@@ 4.2 Sampling Algorithm
In the previous research [35, 26, 47], sample selection is usually modelled by a non-convex combinatorial optimization problem. In contrast, following the idea of [3], we regard $\min_{\mathcal{I}_{m}\in[n]}\operatorname{MMD}^{2}_{k,\alpha}(\mathbf{X}_{\mathcal{I}_{m}},\mathbf{X}_{n})$ as a convex optimization problem by exploiting the convexity of $\alpha$-MMD, and then solve it by a fast iterative minimization procedure derived from Frank-Wolfe algorithm (see Appendix A for derivation details):
@@ 5.1 Generalization Bounds
where $T$ is the full labeled dataset and $S\subset T$ is the core set, $R(h)$ is the expected risk of $h$, $\widehat{R}_{T}(h),\widehat{R}_{S}(h)$ are empirical risk of $h$ on $T,S$. The first term $\widehat{R}_{S}(h)$ is unknown before we label the selected samples, and the second term $|R(h)-\widehat{R}_{T}(h)|$ can be upper bounded by the so-called generalization bounds [2] which do not depend on the choice of core set. Therefore, to control the upper bound of $R(h)$, we only need to analyse the upper bound of the third term $|\widehat{R}_{T}(h)-\widehat{R}_{S}(h)|$ called core-set loss, which requires several mild assumptions.
@@ 7.1 Implementation Details of Our Method
7.2 Comparison with Other Sampling Methods
Main results. We apply RDSS on Flexmatch and Freematch to compare with the following three baselines and two SOTA methods in SSL under different annotation budget settings. The baselines conclude Stratified, Random and $k$-Means, while the two SOTA methods are USL [47] and ActiveFT [54]. The results are shown on Table 1 from which we have several observations: (1) Our proposed RDSS achieves the highest accuracy, outperforming other sampling methods, which underscores the effectiveness of our approach; (2) USL attains suboptimal results under most budget settings yet exhibits a significant gap compared to RDSS, particularly under severely constrained ones. For instance, FreeMatch achieves a $4.95\%$ rise on the STL-10 with a budget of $40$; (3) In most experiments, RDSS either approaches or surpasses the performance of stratified sampling, especially on SVHN and STL-10. However, the stratified sampling method is practically infeasible given that the category labels of the data are not known a priori.
Results on ImageNet. We also compare the second-best method USL with RDSS on ImageNet. Following the settings of FreeMatch [50], we select 100k samples for annotation. FreeMatch, using RDSS and USL as sampling methods, achieves $58.24\%$ and $56.86\%$ accuracy, respectively, demonstrating a substantial enhancement in the performance of our method over the USL approach.
@@ 7.3 Comparison with AL/SSAL Approaches
Second, we compare RDSS with sampling methods used in SSAL when applied to the same SSL framework (i.e., FlexMatch or FreeMatch). The sampling methods conclude CoreSetSSL [35], MMA [39], CBSSAL [11], and TOD-Semi [16]. In detail, we tune recent SSAL approaches with their public implementations and run experiments under an extremely low-budget setting, i.e., 40 samples in a 20-random-and-20-selected setting. Table 3 illustrates that the performance of most SSAL approaches falls below that of random sampling methods under extremely low-budget settings. This inefficiency stems from the dependency of sample selection on model performance within the SSAL framework, which struggles when the model is weak. Our model-free method, in contrast, selects samples before training, avoiding these pitfalls.
@@
Third, when applied to SSL, we directly compare RDSS with the above AL/SSAL approaches, which may better reflect the paradigm differences. The experimental results and analysis are in the Appendix D.6.

### v2

@@ Abstract
Semi-Supervised Learning (SSL) has become a preferred paradigm in many deep learning tasks, which reduces the need for human labor. Previous studies primarily focus on effectively utilising the labelled and unlabeled data to improve performance. However, we observe that how to select samples for labelling also significantly impacts performance, particularly under extremely low-budget settings. The sample selection task in SSL has been under-explored for a long time. To fill in this gap, we propose a Representative and Diverse Sample Selection approach (RDSS). By adopting a modified Frank-Wolfe algorithm to minimise a novel criterion $\alpha$-Maximum Mean Discrepancy ($\alpha$-MMD), RDSS samples a representative and diverse subset for annotation from the unlabeled data. We demonstrate that minimizing $\alpha$-MMD enhances the generalization ability of low-budget learning. Experimental results show that RDSS consistently improves the performance of several popular SSL frameworks and outperforms the state-of-the-art sample selection approaches used in Active Learning (AL) and Semi-Supervised Active Learning (SSAL), even with constrained annotation budgets. Our code is available at RDSS.
@@ 2 Related Work
Semi-Supervised Learning (SSL) effectively utilizes sparse labeled data and abundant unlabeled data for model training. Consistency Regularization [34, 20, 45], Pseudo-Labeling [21, 56] and their hybrid strategies [40, 63, 35] are commonly used in SSL. Consistency Regularization ensures the model’s output stays stable even when there’s noise or small changes in the input, usually from the data augmentation [55]. Pseudo-labelling integrates high-confidence data pseudo-labels directly into training, adhering to entropy minimization [23]. Moreover, an integrative approach that combines the aforementioned strategies can also achieve substantial results [53, 59]. Even though these approaches have been proven effective, they usually assume that labelled samples are randomly selected from each class (i.e., stratified sampling), which is not practical in real-world scenarios where the label for each sample is unknown.
Active Learning
@@
Active learning (AL) aims to optimize the learning process by selecting the appropriate samples for labelling, reducing reliance on large labelled datasets. There are two different criteria for sample selection: uncertainty and representativeness. Uncertainty sampling selects samples about which the current model is most uncertain. Earlier studies utilized posterior probability [22, 49], entropy [18, 26], and classification margin [47] to estimate uncertainty. Recent research regards uncertainty as training loss [17, 60], influence on model performance [11, 24] or the prediction discrepancies between multiple classifiers [8]. However, uncertainty sampling methods may exhibit performance disparities across different models, leading researchers to focus on representativeness sampling, which aims to align the distribution of selected subset with that of the entire dataset [36, 39, 27]. Most AL approaches are difficult to perform well under extremely low-label settings. This may be because they usually require randomly selected samples to begin with and involve human annotators in iterative cycles of labelling and training, leading to substantial labelling overhead.
Model-Free Subsampling
@@
Subsampling is a statistical approach which selects a subset with size $m$ as a surrogate for the full dataset with size $n\gg m$. While model-based subsampling methods depend heavily on the model assumptions [1, 61], improper choice of the model could lead to bad performance of estimation and prediction. In that case, model-free subsampling is preferred in data-driven modelling tasks, as it does not depend on the model assumptions. There are mainly two kinds of popular model-free subsampling methods. The one is induced by minimizing statistical discrepancies, which forces the distribution of subset to be similar to that of full data, in other words, selects representative subsamples, such as Wasserstein distance [13], energy distance [28], uniform design [65], maximum mean discrepancy [7] and generalized empirical $F$-discrepancy [66]. The other tends to select a diverse subset containing as many informative samples as possible [54]. The above-mentioned methodologies either exclusively focus on representativeness or diversity, which are difficult to effectively apply to SSL.
@@ 4 Representative and Diversity Sample Selection
The Representative and Diverse Sample Selection (RDSS) framework consists of two steps: (1) Quantification. We quantify the representativeness and diversity of selected samples by a novel concept called $\alpha$-MMD (6), where $\lambda$ is replaced by $\alpha$ as the trade-off hyperparameter; (2) Optimization. We optimize $\alpha$-MMD by GKHR algorithm to obtain the optimally selected samples $\mathbf{X}_{\mathcal{I}^{*}_{m}}$.
@@ 4.2 Sampling Algorithm
In the previous research [36, 27, 50, 38, 58], sample selection is usually modelled by a non-convex combinatorial optimization problem. In contrast, following the idea of [4], we regard $\min_{\mathcal{I}_{m}\in[n]}\operatorname{MMD}^{2}_{k,\alpha}(\mathbf{X}_{\mathcal{I}_{m}},\mathbf{X}_{n})$ as a convex optimization problem by exploiting the convexity of $\alpha$-MMD, and then solve it by a fast iterative minimization procedure derived from Frank-Wolfe algorithm (see Appendix A for derivation details):
@@ 5.1 Generalization Bounds
where $T$ is the full labeled dataset and $S\subset T$ is the core set, $R(h)$ is the expected risk of $h$, $\widehat{R}_{T}(h),\widehat{R}_{S}(h)$ are empirical risk of $h$ on $T,S$. The first term $\widehat{R}_{S}(h)$ is unknown before we label the selected samples, and the second term $|R(h)-\widehat{R}_{T}(h)|$ can be upper bounded by the so-called generalization bounds [3, 64] which do not depend on the choice of core set. Therefore, to control the upper bound of $R(h)$, we only need to analyse the upper bound of the third term $|\widehat{R}_{T}(h)-\widehat{R}_{S}(h)|$ called core-set loss, which requires several mild assumptions. Shalit, et al. [37] derived a MMD-type upper bound for $|\widehat{R}_{T}(h)-\widehat{R}_{S}(h)|$ to estimate individual treatment effect, while our bound is generalized to a wider range of tasks.
@@ 7.1 Implementation Details of Our Method
7.2 Comparison with Other Sampling Methods
Main results We apply RDSS on Flexmatch and Freematch to compare with the following three baselines and two SOTA methods in SSL under different annotation budget settings. The baselines conclude Stratified, Random and $k$-Means, while the two SOTA methods are USL [50] and ActiveFT [57]. The results are shown on Table 1 from which we have several observations: (1) Our proposed RDSS achieves the highest accuracy, outperforming other sampling methods, which underscores the effectiveness of our approach; (2) USL attains suboptimal results under most budget settings yet exhibits a significant gap compared to RDSS, particularly under severely constrained ones. For instance, FreeMatch achieves a $4.95\%$ rise on the STL-10 with a budget of $40$; (3) In most experiments, RDSS either approaches or surpasses the performance of stratified sampling, especially on SVHN and STL-10. However, the stratified sampling method is practically infeasible given that the category labels of the data are not known a priori.
Results on ImageNet We also compare the second-best method USL with RDSS on ImageNet. Following the settings of FreeMatch [53], we select 100k samples for annotation. FreeMatch, using RDSS and USL as sampling methods, achieves $58.24\%$ and $56.86\%$ accuracy, respectively, demonstrating a substantial enhancement in the performance of our method over the USL approach.
@@ 7.3 Comparison with AL/SSAL Approaches
Second, we compare RDSS with sampling methods used in SSAL when applied to the same SSL framework (i.e., FlexMatch or FreeMatch) on CIFAR-10. The sampling methods conclude CoreSetSSL [36], MMA [42], CBSSAL [12], and TOD-Semi [17]. In detail, we tune recent SSAL approaches with their public implementations and run experiments under an extremely low-budget setting, i.e., 40 samples in a 20-random-and-20-selected setting. Table 3 illustrates that the performance of most SSAL approaches falls below that of random sampling methods under extremely low-budget settings. This inefficiency stems from the dependency of sample selection on model performance within the SSAL framework, which struggles when the model is weak. Our model-free method, in contrast, selects samples before training, avoiding these pitfalls.
@@
Third, we directly compare RDSS with the above AL/SSAL approaches when applied to SSL, which may better reflect the paradigm differences. The experimental results and analysis are in the Appendix D.6.
@@ 8 Conclusion
Acknowledgements
This research was partially supported by National Natural Science Foundation of China under grant No. 82202984, Zhejiang Key R&\&&D Program of China under grants No. 2023C03053 and No. 2024SSYS0026, and US National Science Foundation under grant No. 2316011.
@@ Appendix C Proof of Theorems
The proof borrows the technique introduced in [37] for decomposing the expected risk of hypotheses.